Annual starting salaries for college graduates with degrees in business administration are generally expected to be between $10,000 and $35,000. Assume that a 95% confidence interval estimate of the population mean annual starting salary is desired.

a. What is the planning value for the population standard deviation?

b. How large a sample should be taken if the desired margin of error is $400? Round your answer to next whole number.

$230?

$90?
Answer #1

a)

std dev,σ = Range/4 = (35000-10000)/4 = 6250

b)

i) Standard Deviation ,   σ =    6250                  
sampling error ,    E =   400                  
Confidence Level ,   CL=   95%                  
                          
alpha =   1-CL =   5%                  
Z value =    Zα/2 =    1.960   [excel formula =normsinv(α/2)]              
                          
Sample Size,n = (Z * σ / E )² = (   1.960   *   6250   /   400   ) ² =   937.856
                          
                          
So,Sample Size needed=       938                  

---------------

ii)

Standard Deviation ,   σ =    6250                  
sampling error ,    E =   230                  
Confidence Level ,   CL=   95%                  
                          
alpha =   1-CL =   5%                  
Z value =    Zα/2 =    1.960   [excel formula =normsinv(α/2)]              
                          
Sample Size,n = (Z * σ / E )² = (   1.960   *   6250   /   230   ) ² =   2836.616
                          
                          
So,Sample Size needed=       2837                  
------------------

iii)

Standard Deviation ,   σ =    6250                  
sampling error ,    E =   90                  
Confidence Level ,   CL=   95%                  
                          
alpha =   1-CL =   5%                  
Z value =    Zα/2 =    1.960   [excel formula =normsinv(α/2)]              
                          
Sample Size,n = (Z * σ / E )² = (   1.960   *   6250   /   90   ) ² =   18525.554
                          
                          
So,Sample Size needed=       18526